# Networked Communication for Decentralised Cooperative Agents in Mean-Field Control

### v1

@@ 2.2 Munchausen Online Mirror Descent
Recent works have solved MFGs from non-episodic runs of the finite empirical system using a form of policy iteration called Online Mirror Descent (OMD) (Benjamin & Abate, 2024); we adapt this to learn a social optimum in the MFC setting. OMD involves beginning with an initial policy $\pi_{0}$, and then at each iteration $k$, evaluating the current policy $\pi_{k}$ with respect to its induced mean-field flow $\boldsymbol{\mu}=I(\pi_{k})$ to compute its Q-function $Q_{k+1}$. To stabilise the learning process, we then use a weighted sum over this and past Q-functions, and set $\pi_{k+1}$ to be the softmax over this weighted sum, i.e. $\pi_{k+1}(\cdot|o)=softmax\left(\frac{1}{\tau_{q}}\sum^{k+1}_{\kappa=0}{Q_{\kappa}}(o,\cdot)\right)$. $\tau_{q}$ is a temperature parameter that scales the entropy in Munchausen RL (Vieillard et al., 2020); this is a different temperature to the one agents use when communicating policies, denoted $\tau^{comm}_{k}$ and discussed in Sec. 3.3.
If the Q-function is approximated non-linearly, it is difficult to compute this weighted sum. The Munchausen trick addresses this by computing a single Q-function that mimics the weighted sum using implicit regularisation based on the Kullback-Leibler (KL) divergence between $\pi_{k}$ and $\pi_{k+1}$ (Vieillard et al., 2020). Using this reparametrisation gives Munchausen OMD (MOMD), detailed in Sec. 3.2 (Laurière et al., 2022b; Wu et al., 2024). MOMD does not bias policies, and has the same convergence guarantees as OMD (Hadikhanloo, 2017; Perolat et al., 2021; Wu et al., 2024).
@@ 2.3 Networks
The decentralised population has the following communication graph:
Definition 7 (Time-varying communication network).
[…]
@@ 3 Learning and estimation algorithms
We adapt recent algorithms for the MFG setting, where networked communication is used 1) to form local estimates of the global empirical mean field, and 2) to allow agents to adopt better-performing policy updates from their neighbours to accelerate learning (Benjamin & Abate, 2024). We adapt these algorithms for cooperative MFC, where decentralised agents must now optimise the population-average return instead of their individual one (the decentralised agents may not always follow a common policy while training unless we make strong assumptions on the communication network as in Sec. 4, so we do not directly optimise the social welfare from Def. 5).
It is unrealistic to assume that decentralised agents would have access to the global average reward, so we identify a third use of the communication network, namely 3) to allow agents to estimate the global average reward $\hat{r}_{t}$ from a local neighbourhood. We contribute a novel algorithm Alg. 1 for this purpose described in Sec. 3.1, and we give our main learning method Alg. 2 in Sec. 3.2. Meanwhile Alg. 3 for estimating the mean field, which is taken from Alg. 3 in Benjamin & Abate (2024) for the MFG setting, is described in Appx. B.1. Our policy communication algorithm Alg. 4 is also based on that in Benjamin & Abate (2024) for the MFG setting and is therefore placed in Appx. B, but since it is key to our novel theoretical results that we contribute for the MFC setting, we give a description of Alg. 4 in the main text in Sec. 3.3.
3.1 Sub-routine for networked estimation of global average reward
@@
1:  $\forall i:$ Initialise reward sets ^$\hat{\mathcal{R}}^{i}_{t}\leftarrow\{(ID^{i},r^{i}_{t})\}$
@@
3:     $\forall i:$ Broadcast ^$\hat{\mathcal{R}}^{i}_{t,c_{r}}$
@@
7:  $\forall i:$ $\tilde{\hat{r}}^{i}_{t}\leftarrow\frac{1}{|\hat{\mathcal{R}}^{i}_{t,C_{r}}|}\sum_{(ID,r)\in\hat{\mathcal{R}}^{i}_{t,C_{r}}}r$
@@
Our novel Alg. 1 involves agents using the communication network $\mathcal{G}^{comm}_{t}$ to locally estimate the global population-average reward received after a given step in the environment - maximising the average reward ensures agents are solving the cooperative MFC problem instead of the non-cooperative MFG. Agents broadcast their received reward with a unique ID to ensure each reward is only counted once (Line 1). They collect those received from neighbours, and repeat the process of broadcasting and expanding their collections for a further $C_{r}-1$ rounds, so as to receive rewards from agents more than one hop away on the network (Lines 2-6). They set their estimate of the global average to the average of the rewards they have collected within the rounds (Line 7).
3.2 Main algorithm for updating Q-networks and policies
@@
1:  $\forall i$ : Randomly initialise parameters $\theta^{i}_{0}$ of Q-networks $\check{Q}_{\theta^{i}_{0}}(o,\cdot)$, and set $\pi^{i}_{0}(a|o)=$ softmax$\left(\frac{1}{\tau_{q}}\check{Q}_{\theta^{i}_{0}}(o,\cdot)\right)(a)$
2:  for $k=0,\dots,K-1$ do
@@
4:     for $m=0,\dots,M-1$ do
5:        $\{o^{i}_{t}\}_{i=1}^{N}\leftarrow$ EstimateMeanFieldAlg. 3$\left(\mathcal{G}^{vis}_{t},\mathcal{G}^{comm}_{t},\{s^{i}_{t}\}_{i=1}^{N}\right)$
@@
7:        $\{\tilde{\hat{r}}^{i}_{t}\}_{i=1}^{N}\leftarrow$ EstimateAverageRewardAlg. 1$\left(\mathcal{G}^{comm}_{t},\{r^{i}_{t}\}_{i=1}^{N}\right)$
@@
12:        Update $\theta$ to minimise ^$\hat{\mathcal{L}}(\theta,\theta^{\prime})$ as in Def. 8
@@
16:     $\forall i$ : $\pi^{i}_{k+1}(a|o)\leftarrow$ softmax$\left(\frac{1}{\tau_{q}}\check{Q}_{\theta^{i}_{k+1}}(o,\cdot)\right)(a)$
17:     ($\left(\{\pi^{i}_{k+1}\}_{i},\{s^{i}_{t}\}_{i},t\right)$leftarrow CommunicatePolicyAlg. 4$\left(\mathcal{G}^{comm}_{t},\{\pi^{i}_{k+1}\}_{i},\{s^{i}_{t}\}_{i},t\right)$
@@
Our novel Alg. 2, adapted from non-cooperative Alg. 1 in Benjamin & Abate (2024), contains the core method for online MFC learning using the empirical mean field in a non-episodic system run. Each agent $i$ approximates its Q-function $\check{Q}_{\theta^{i}_{k}}(o,\cdot)$ with its own neural network parametrised by $\theta^{i}_{k}$. Agent $i$’s policy is determined by $\pi_{\theta^{i}_{k}}(a|o)=$ softmax$\left(\frac{1}{\tau_{q}}\check{Q}_{\theta^{i}_{k}}(o,\cdot)\right)(a)$ we denote this as $\pi^{i}_{k}(a|o)$ for simplicity when appropriate. Each agent maintains a buffer (with size $M$) of collected transitions of the form $\left(o^{i}_{t},a^{i}_{t},\tilde{\hat{r}}^{i}_{t},o^{i}_{t+1}\right)$, where $\tilde{\hat{r}}^{i}_{t}$ is agent $i$’s local estimate of the global average reward obtained by running Alg. 1 (Line 7).
At each iteration $k$, agents empty their buffer (Line 3) before collecting $M$ new transitions in the environment (Lines 4-9). Each decentralised agent then trains its Q-network $\check{Q}_{\theta^{i}_{k}}$ via $L$ training updates (Lines 10-14) as follows. For training purposes, $i$ also maintains a target network $\check{Q}_{\theta^{i,^{\prime}}_{k,l}}$ with the same architecture but parameters $\theta^{i,^{\prime}}_{k,l}$ copied from $\theta^{i}_{k,l}$ less regularly than $\theta^{i}_{k,l}$ themselves are updated, i.e. only every $\nu$ learning iterations (Line 13). At each iteration $l$, the agent samples a random batch $B^{i}_{k,l}$ of $|B|$ transitions from its buffer (Line 11). It then trains its neural network using stochastic gradient descent to minimise the following empirical loss (Line 12):
Definition 8 (Empirical loss for Q-network).
The empirical loss is given by
^$\hat{\mathcal{L}}(\theta,\theta^{\prime})=\frac{1}{|B|}\sum_{transition\in B^{i}_{k,l}}\left|\check{Q}_{\theta^{i}_{k,l}}(o_{t},a_{t})-T\right|^{2},\quad\text{where the target $T$ is}$
$T=\tilde{\hat{r}}_{t}\;+\;\left[\tau_{q}\ln\pi_{\theta^{i,^{\prime}}_{k,l}}(a_{t}|o_{t})\right]^{0}_{cl}+\gamma\sum_{a\in\mathcal{A}}\pi_{\theta^{i,^{\prime}}_{k,l}}(a|o_{t+1})\left(\check{Q}_{\theta^{i,^{\prime}}_{k,l}}(o_{t+1},a)-\tau_{q}\ln\pi_{\theta^{i,^{\prime}}_{k,l}}(a|o_{t+1})\right).$
For $cl<0$, $[\cdot]^{0}_{cl}$ is a clipping function used in Munchausen RL to prevent numerical issues if the policy is too close to deterministic, as the log-policy term is otherwise unbounded (Vieillard et al., 2020; Wu et al., 2024). The newly trained Q-network determines $i$’s updated policy (Line 16).
@@ 3.3 Sub-routine for communicating and refining policies
Alg. 4 (Appx. B.2, based on Alg. 1 in Benjamin & Abate (2024) for the MFG setting) uses the communication network $\mathcal{G}^{comm}_{t}$ to spread policy updates that are estimated to be better performing through the population, allowing faster learning than in the independent and even centralised cases.
@@
Alg. 4 is run after agents have independently updated their policies according to their newly trained Q-networks at each iteration $k$ of the main learning algorithm (Line 17, Alg. 2). In Alg. 4, agents obtain an approximation of their individual discounted expected return $\{{V}^{i}(\boldsymbol{\pi},\mu_{0})\}_{i=1}^{N}$ (Def. 2, i.e. not the population-average return, which would not give differentiation between the different updated policies). They do so by collecting individual rewards for $E$ steps, and calculating the discounted sum of rewards over these finite steps, setting this value to $\sigma^{i}_{k+1}$ (Lines 1-7). We can characterise this approximation of the infinite-step return as $\sigma^{i}_{k+1}$}${}_{i=1}^{N}$ = $\{\hat{V}^{i}(\boldsymbol{\pi}_{k+1},\mu_{0};E)\}_{i=1}^{N}$.
@@
They then broadcast their Q-network parameters along with $\sigma^{i}_{k+1}$ (Line 9). Receiving these from their neighbours $J^{i}_{t}$ on the communication network, agents select which set of parameters to adopt by taking a softmax over their own and the received estimate values $\sigma^{j}_{k+1}$ $\forall j\in J^{i}_{t}$, defined as follows: $\mathrm{adopted}^{i}\sim$ Pr$\left(\mathrm{adopted}^{i}=j\right)$ = $\frac{\exp{(\sigma^{j}_{k+1}}/\tau^{comm}_{k})}{\sum_{x\in J^{i}_{t}}\exp{(\sigma^{x}_{k+1}}/\tau^{comm}_{k})}$ (Lines 10-12). They can repeat this broadcast and adoption process for $C_{p}$ rounds (distinct from the $C_{r}$ and $C_{e}$ communication rounds for the other sub-routines). We theoretically prove the benefits of this method in the following.

### v2

@@ 3.2 Munchausen Online Mirror Descent
Recent works have solved MFGs from non-episodic runs of the finite-population empirical system using a form of policy iteration called Online Mirror Descent (OMD) (benjamin2024networkedapproximation); we adapt this to learn a social optimum in the MFC setting. OMD involves beginning with an initial policy $\pi_{0}$, and then at each iteration $k$, evaluating the current policy $\pi_{k}$ with respect to its induced mean-field flow $\bm{\mu}=I(\pi_{k})$ to compute its Q-function $Q_{k+1}$. To stabilise the learning process, we then use a weighted sum over this and past Q-functions, and set $\pi_{k+1}$ to be the softmax over this weighted sum, i.e. $\pi_{k+1}(\cdot|o)=softmax\left(\frac{1}{\tau_{q}}\sum^{k+1}_{\kappa=0}{Q_{\kappa}}(o,\cdot)\right)$. $\tau_{q}$ is a temperature parameter that scales the entropy in Munchausen RL (see Sec. 4.2) (NEURIPS2020_2c6a0bae); this is a different temperature to the one agents use when communicating policies, denoted $\tau^{comm}_{k}$ and discussed in Sec. 4.3.
If the Q-function is approximated non-linearly, it is difficult to compute this weighted sum. The Munchausen trick addresses this by computing a single Q-function that mimics the weighted sum using implicit regularisation based on the Kullback-Leibler (KL) divergence between $\pi_{k}$ and $\pi_{k+1}$ (NEURIPS2020_2c6a0bae). Using this reparametrisation gives Munchausen OMD (MOMD), detailed in Sec. 4.2 (scalable_deep; wu2024populationaware). MOMD does not bias policies, and has the same convergence guarantees as OMD (hadikhanloo2017learninganonymousnonatomicgames; perolat2021scaling; wu2024populationaware).
@@ 3.3 Networks
Our decentralised population exhibits two time-varying graphs. The first is a communication network, by which agents can exchange information:
Definition 3.9 (Time-varying communication network).
The time-varying graph ($\mathcal{G}^{comm}_{t}$)t≥0 is given by $\mathcal{G}^{comm}_{t}$ = ($\mathcal{N},\mathcal{E}^{comm}_{t}$), where $\mathcal{N}$ is the set of vertices each representing an agent $i\in\{1,\dots,N\}$, and the edge set $\mathcal{E}_{t}$ $\subseteq$ {(i,j) : i,j $\in$ $\mathcal{N}$, i $\neq$ j} is the set of undirected links present at time t. A network’s diameter $d_{\mathcal{G}^{comm}_{t}}$ is the maximum of the shortest path lengths between any pair of nodes.
In principle, agents can use this same communication network to receive information about others’ state in order to estimate the mean field. However, we also define an alternative observation graph that is useful in a specific subclass of environments, which can most intuitively be thought of as those where agents’ states are positions in physical space, which include those in our experiments. When this is the case, we usually think of agents’ ability to observe each other as depending more abstractly on whether states are visible to each other. This visibility graph is:
Definition 3.10 (Time-varying state-visibility graph).
The time-varying state visibility graph ($\mathcal{G}^{vis}_{t}$)t≥0 is given by $\mathcal{G}^{vis}_{t}$ = ($\mathcal{S^{\prime}},\mathcal{E}^{vis}_{t}$), where $\mathcal{S^{\prime}}$ is the set of vertices representing the environment states $\mathcal{S}$, and the edge set $\mathcal{E}^{vis}_{t}$ $\subseteq$ {(m,n) : m,n $\in$ $\mathcal{S^{\prime}}$} is the set of undirected links present at time t, indicating which states are visible to each other.
In Sec. 4.4 we present Alg. 4, which forms an initial estimate of the global empirical mean field (to serve as an observation input for agents’ Q-/policy-networks) via the visibility graph $\mathcal{G}^{vis}_{t}$, before refining this estimate via the communication graph $\mathcal{G}^{comm}_{t}$. benjamin2024networkedapproximation discusses an algorithm for more general settings where the visibility graph $\mathcal{G}^{vis}_{t}$ does not apply.
@@ 4 Learning and estimation algorithms
We adapt recent algorithms for the MFG setting, where networked communication is used 1) to form local estimates of the global empirical mean field, and 2) to allow agents to adopt better-performing policies from neighbours to accelerate learning (benjamin2024networkedapproximation). We adapt these algorithms for cooperative MFC, where decentralised agents must optimise the population-average return instead of their individual one (the decentralised agents may not always follow a common policy while training unless we make strong assumptions on the communication network as in Sec. 5, so we do not directly optimise social welfare from Def. 3.5).
It is unrealistic to assume that decentralised agents have access to the global average reward, so we find a third use of the communication network in 3) allowing agents to estimate the global average reward $\hat{r}_{t}$ from a local neighbourhood. We contribute a novel algorithm Alg. 1 for this purpose (Sec. 4.1), and we describe our main learning method Alg. 2 in Sec. 4.2. Our policy communication algorithm Alg. 3, based on that in benjamin2024networkedapproximation for the MFG setting, is described in Sec. 4.3. Meanwhile Alg. 4 for estimating the mean field, which is taken from Alg. 3 in benjamin2024networkedapproximation for the MFG setting, is described in Sec. 4.4.
@@
[…]
@@
3:  $\forall i:$ Broadcast $\hat{\mathcal{R}}^{i}_{t,c_{r}}$
@@
[…]
@@
4.1 Sub-routine for networked estimation of global average reward
Our novel Alg. 1 involves agents using the communication network $\mathcal{G}^{comm}_{t}$ to locally estimate the global population-average reward received after a given step in the environment. Maximising the population-average reward ensures agents are solving the cooperative MFC problem instead of the non-cooperative MFG. Agents broadcast their received reward with a unique ID to ensure each reward is only counted once (Line 1). They collect those received from neighbours, and repeat the process of broadcasting and expanding their collections for a further $C_{r}-1$ rounds, so as to receive rewards from agents more than one hop away on the network (Lines 2-6). They finally set their estimate of the global average to the average of the rewards they have collected (Line 7).
4.2 Main learning algorithm for updating Q-networks and policies
@@
[…]
2:for $k\in 0,\dots,K-1$ do
@@
4:  for $m\in 0,\dots,M-1$ do
[…]
@@
[…]
@@
12:   Update $\theta$ to minimise $\hat{\mathcal{L}}(\theta,\theta^{\prime})$ as in Def. 4.1
@@
[…]
@@
Our novel Alg. 2, adapted from non-cooperative Alg. 1 in benjamin2024networkedapproximation, contains the core method for online MFC learning using the empirical mean field in a non-episodic system run. Our MOMD-based method (Sec. 3.2) works as follows. Each agent $i$ approximates its Q-function $\check{Q}_{\theta^{i}_{k}}(o,\cdot)$ with its own neural network parametrised by $\theta^{i}_{k}$. Agent $i$’s policy is determined by
$\pi_{\theta^{i}_{k}}(a|o)=\text{softmax}\left(\frac{1}{\tau_{q}}\check{Q}_{\theta^{i}_{k}}(o,\cdot)\right)\left(a\right).$
We denote this as $\pi^{i}_{k}(a|o)$ for simplicity when appropriate. Each agent maintains a buffer (with size $M$) of collected transitions of the form $\left(o^{i}_{t},a^{i}_{t},\tilde{\hat{r}}^{i}_{t},o^{i}_{t+1}\right)$, where $\tilde{\hat{r}}^{i}_{t}$ is $i$’s local estimate of the global average reward obtained by running Alg. 1 (Line 7). At each iteration $k$, agents empty their buffer (Line 3) before collecting $M$ new transitions in the environment (Lines 4-9). Each decentralised agent then trains its Q-network $\check{Q}_{\theta^{i}_{k}}$ via $L$ updates (Lines 10-14) as follows.
For stability, $i$ also maintains a target network $\check{Q}_{\theta^{i,^{\prime}}_{k,l}}$ with the same architecture but parameters $\theta^{i,^{\prime}}_{k,l}$ copied from $\theta^{i}_{k,l}$ less regularly than $\theta^{i}_{k,l}$ themselves are updated, i.e. only every $\nu$ learning iterations (Line 13). At each iteration $l$, the agent samples a random batch $B^{i}_{k,l}$ of $|B|$ transitions from its buffer (Line 11). It then trains its Q-network using stochastic gradient descent to minimise the loss in Def 4.1 (Line 12). The trained Q-network determines $i$’s updated policy (Line 16).
Definition 4.1 (Q-network empirical loss).
The training loss to be minimised is given by
$\hat{\mathcal{L}}(\theta,\theta^{\prime})=\frac{1}{|B|}\sum_{transition\in B^{i}_{k,l}}\left|\check{Q}_{\theta^{i}_{k,l}}(o_{t},a_{t})-T\right|^{2},$
$\text{where}\;\;\;\;T=\tilde{\hat{r}}_{t}+\left[\tau_{q}\ln\pi_{\theta^{i,^{\prime}}_{k,l}}(a_{t}|o_{t})\right]^{0}_{cl}+\gamma\sum_{a\in\mathcal{A}}\pi_{\theta^{i,^{\prime}}_{k,l}}(a|o_{t+1})\left(\check{Q}_{\theta^{i,^{\prime}}_{k,l}}(o_{t+1},a)-\tau_{q}\ln\pi_{\theta^{i,^{\prime}}_{k,l}}(a|o_{t+1})\right).$
For $cl<0$, $\cdot]^{0}_{cl}$ is a clipping function used in Munchausen RL to prevent numerical issues if the policy is too close to deterministic, as the log-policy term is otherwise unbounded (NEURIPS2020_2c6a0bae; wu2024populationaware).
@@ 4.3 Sub-routine for communicating and refining policies
Algorithm 3 Policy communication and selection
0: policies $\pi^{i}_{k+1}$$_{i=1}^{N}$; states $s^{i}_{t}$$_{i=1}^{N}$; $t$
[…]
2:for $e\in 0,\dots,E-1$ evaluation steps do
[…]
6:  $t\leftarrow t+1$
7:end for
8:for $C_{p}$ rounds do
9:  $\forall i:$ Broadcast $\sigma^{i}_{k+1},\pi^{i}_{k+1}$
10:  $\forall i:J^{i}_{t}\leftarrow i\cup\{j\in\mathcal{N}:(i,j)\in\mathcal{E}^{comm}_{t}$}
[…]
15:end for
16:return (policies $\pi^{i}_{k+1}$$_{i=1}^{N}$, states $s^{i}_{t}$$_{i=1}^{N}$, $t$)
@@
Alg. 3 (based on Alg. 1 in benjamin2024networkedapproximation for MFGs) uses the communication network $\mathcal{G}^{comm}_{t}$ to spread policy updates that are estimated to be better performing through the population, allowing faster learning than in the independent and central-agent cases.
@@
Alg. 3 is run after agents have independently updated their policies according to their newly trained Q-networks at each iteration $k$ of the main learning algorithm (Line 17, Alg. 2). In Alg. 3, agents obtain an approximation of their individual discounted expected return $\{{V}^{i}(\bm{\pi},\mu_{t})\}_{i=1}^{N}$ (Def. 3.2, i.e. not the population-average return, which would not give differentiation between the different updated policies). They do so by collecting individual rewards for $E$ steps (not added to the training buffer), and calculating the discounted sum of rewards over these finite steps, setting this value to $\sigma^{i}_{k+1}$ (Lines 1-7). We can characterise this approximation of the infinite-step return as $\sigma^{i}_{k+1}$$_{i=1}^{N}$ = $\{\hat{V}^{i}(\bm{\pi}_{k+1},\mu_{t};E)\}_{i=1}^{N}$.
@@
They then broadcast their Q-network parameters along with $\sigma^{i}_{k+1}$ (Line 9). Receiving these from their neighbours $J^{i}_{t}$ on the network, agents select which set of parameters to adopt by taking a softmax over their own and the received estimate values $\sigma^{j}_{k+1}$ $\forall j\in J^{i}_{t}$, defined as follows (Lines 10-12):
$\mathrm{adopted}^{i}\sim\text{Pr}\left(\mathrm{adopted}^{i}=j\right)=\frac{\exp{(\sigma^{j}_{k+1}}/\tau^{comm}_{k})}{\sum_{x\in J^{i}_{t}}\exp{(\sigma^{x}_{k+1}}/\tau^{comm}_{k})}.$
They repeat this broadcast and adoption process for $C_{p}$ rounds (distinct from the $C_{r}$$C_{e}$ communication rounds for the other sub-routines).
4.4 Sub-routine for networked estimation of global empirical mean-field
Networked agents use Alg. 4 (this is Alg. 3 from benjamin2024networkedapproximation for the MFG setting) to locally estimate the global empirical mean field, to serve as an observation input for their Q-/policy-networks. Recall that we include this added observation and sub-routine for generality, especially for non-stationary problems. However it is often not necessary, particularly in stationary problems like those in our experiments, where agents can find the social optimum while only observing $o^{i}_{t}=s^{i}_{t}$, and therefore would not need to estimate the mean field.
Alg. 4 involves agents using the visibility graph $\mathcal{G}^{vis}_{t}$ to count the number of agents in locations that fall within the visibility radius (Line 2). For $C_{e}$ communication rounds, agents can supplement this local count with those received from neighbours over the communication network $\mathcal{G}^{comm}_{t}$, in order to count agents that do not fall within the visibility radius (Lines 3-8). We assume agents know the population’s total size $N$, and therefore can distribute the uncounted agents uniformly over the states that remain unaccounted for after the communication rounds (Lines 9-11). Agents now have a vector containing a true or estimated count for every state; this is converted to an estimated empirical mean field by dividing all counts by $N$ (Lines 12-13).
[…]
3:for $c_{e}$ $\in$ $1,\dots,C_{e}$ do
4:  $\forall i:$ Broadcast $\hat{\upsilon}^{i}_{t,c_{e}}$
5:  $\forall i:J^{i}_{t}\leftarrow i\cup\{j\in\mathcal{N}:(i,j)\in\mathcal{E}^{comm}_{t}\}$
6:  $\forall i,s:$ Initialise new count vector $\hat{\upsilon}^{i}_{t,(c_{e}+1)}[s]$ with $\emptyset$
7:  $\forall i,s$ and $\forall j\in J^{i}_{t}:\hat{\upsilon}^{i}_{t,(c_{e}+1)}[s]\leftarrow\hat{\upsilon}^{j}_{t,c_{e}}[s]$ if $\hat{\upsilon}^{j}_{t,c_{e}}[s]\neq\emptyset$
8:end for
[…]